• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 16:03:13 on December21,2009

    Entry number 308723

    keyword=Day shift summary

    Day shift summary:

    shift workers: J. P. Chen (SL), R. Michaels (TO), D. Armstrong (3rd)

    8:00 Shift started with target warm due to earlier cryo flow dropped
    (~7:00 am). Started to get cryo flow, in the process of cooling-down.
    9:15 target ready. Requested to MCC to mask target FSD to move to LD2,
    but MCC is doing emittance measurements, told us to wait.
    9:35 Moved target to LD2.
    10:00 MCC finished emittance test. asked for 2 uA for spot++.
    10:05 take runf or spot++ and with VDC on, Q2 measurement.
    10:15 turned off VDC, asked for high current for production. Got 115 uA.
    Smooth production since then.
    12:15 HV alarm: Compton HVs 7, 8. Also noticed Ch. 1 was off. Talked to Alexandre Camsonne, he informed me that he turned off ch. 1 because
    high back ground. Compton is not taking data. He suggested to also turn off HV channels 7, 8 (Compton diagonistics), which I did.
    12:15 QAD crashed, restarted.


    run information

    HAPPEX runs
    14808 10:16-11:14, 100K, 115 uA on LD2
    14809 11:14-12:10, 100K
    14810 12:15-13:12, 100K
    14811 13:13-14:08, 100K
    14813 14:09-15:06, 100K
    14814 15:07-, still on-going at the end of shift.

    HRS runs L/R
    26419/5939 short, PS not set
    26420/5940, 10:09-10:12, 593K/527K, Q2&spot++, 2uA on LD2, VDC on
    26421/5941, 10:18-11:12, 5.56M/5.28M, production, 115uA on LD2, VDC off
    26422/5942, 11:15-12:16, 6.67M/6.53M
    26423/5943, 12:16-13:13, 6.11M/5.84M
    26424/5944, 13:12-14:13, 6.41M/6.16M
    26425/5945, 14:13-15:07, 5.4M/5.16M
    26426/5946, 15:08-, still on-going at the end of shift.